OpenStudy (anonymous):

linear approximation of sqrt(82)

OpenStudy (anonymous):

I understand the concept of linearization and linear approximation. However unsure of calculating a simple single term such as the above. Would I do the following: Using the linearization law of that: f(x) ~~ l(x) = f(a) - f'(a)(x-a) In the case above we would LET a = 80. Thus f(x) = sqrt(x)

OpenStudy (phi):

you use the first 2 terms of a taylor series expansion http://en.wikipedia.org/wiki/Taylor_series#Definition f(x) \(\approx\)f(a) + f'(a) (x-a)/1! in your case, you would pick the nearest number that has an integer square root (rather than an irrational number) i.e. pick a=81, so f(a)=9 \[ f'(x) = \frac{1}{2}x^{-\frac{1}{2}} \] evaluated at x=81, you get f'(a)= 1/18 your approximation is \[f(x) ≈ 9 + \frac{x-81}{18} \] or \[f(x) ≈ 4.5 + \frac{x}{18} \]